OpenStudy (anonymous):

please help Which expression is equivalent to 12a + 8? 4(3a + 2) 4(3a + 8) 6(6a + 2) 6(6a + 8)

OpenStudy (anonymous):

whats ccommon is 4 so you can figure out the ans

OpenStudy (some.random.cool.kid):

this is splitting the 12 with distributive property so probably 6(6 but you have to figure out whether its c or D k :}

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

what is the gcf of 12a + 8

OpenStudy (campbell_st):

what if I wrote both terms as the product of there factors \[4 \times 3 \times a + 4 \times 2\] take the common factor and write it outside the brackets what's left is inside the brackets

OpenStudy (anonymous):

I think its the first one because 4 x 3 = 12 and 4 x 2 = 8 so it would equal 12a + 8